OpenStudy (anonymous):

A girl has 8 coins in dimes and quarters. Their value is $1.55. How many quarters does she have?

OpenStudy (mertsj):

d+q=8 .10d+.25q=1.55

OpenStudy (lollylau):

Diagram method. 1. Draw 8 circles, representing 8 coins A coin must at least carry a value of 10 cents. 2. Draw 1 leg for 10 cents onto every circle. Draw the remaining legs until the value reaches $1.55. (3 legs for a quarter)

OpenStudy (mertsj):

d=8-q from the first equation. Replace d with 8-q in the second equation and solve.

OpenStudy (lollylau):

I still like my method better :)

OpenStudy (mertsj):

.10(8-q)+.25q=1.55 10(8-q)+25q=155 80-10q+25q=155 15q=75 q=5 d=3

OpenStudy (anonymous):

thanks alot

OpenStudy (lollylau):

your welcome :)